Last of Berry's updates from the field

Greetings again from Crary Lab on a Sunday afternoon. This week has been full of starts and stops with interruptions due to weather. Difficulties with the weather, and hence flying, have led to a less productive week in the field than the past two.
504- Anna continues to work out of Lake Hoare Camp. She is sampling melt water on and from the Canada Glacier and also did an albedo measurement this week. Adam has helped both Christa and Aneliya with their work, and he is currently conducting a hyporheic zone experiment at F6 with the help of my 2 students, Sydney and Elsa. 
505- The limno team finished the limno 'lightning' round yesterday at Miers. Noelle and Jeff head North tomorrow, while Maddie will stay on a few more days to help Amy. Amy has been extremely busy doing CO2 and chl-a measurements from the lakes. The part for the sed-trap came in yesterday, so Amy and Maddie will redeploy the trap and finished some bioassays with moat water this coming week.
506- Christa ran a tracer experiment at Priscu stream and hopes to get a couple more in, but the weather (cold, windy, overcast) has greatly affected stream flow. Mikey, Zach and Forrest have been doing their stream work and helping out Christa, and Aneliya as noted above. Aneliya began 'algal ops' in the variousTV streams and I hear all goes well. 
507- Jeb arrived earlier this week, and Ashley arrived last night so the Wormherders are at full strength. They had a very successful first week in that they sampled the P3 experiment, as well as the plots on the Southside of LH, plus Byron, Jeb, and Dan are at F6 and began the manipulations at P3 yesterday. I think that they are very pleased at their progress. I know that both Byron and Jeb have been working on MCM V writing assignments too. 
509- Although bad weather prevented a second trip to Miers Valleys all the other stream work dealing with determining chemical fluxes into the ocean was completed. As noted above Syd and Elsa are out helping Adam as well as doing a diurnal sampling of VG for their own work. Kathy is completely caught up with the core LTER ion analyses, while she and Elsa have analyzed the DOCs through Limno run #2. The new TOC instrument is now up and running and the first batch of standards look very good. Elsa will start running stream samples with it this week.
Add-on-Our great colleague to the LTER, Joe Levy, arrived yesterday with 3 young associates, and, after training, they hope to get out, first to TV, by the end of the week.
The weather! Hopefully this week will be sunny and warm. The cargo deadline is closing in, and there'll be much activity in packing, etc this week.
This will be my last report as POC as I am scheduled to head N tomorrow. Although the weather is not looking good, I assume that I will be in Columbus prior to the due date of the next report!! So I will pass the POC torch to our distinguished colleague, Prof Wall. Before I sign off for the last time, I would be remiss without acknowledging PHI, help ops, Liz K and the Crary Lab staff for their great support and professionalism. Finally, thank you, thank you to Rae and Renee for all they do for us and especially the young folks on our MCM-LTER team. I wish all still here a great, and successful finish to the season. I wish my young (and old--DM and JP!!$$) colleagues, inspired writing on the proposal-V; may the Muses be with you!

Article Section: